Slick Corporation is a small producer of synthetic motor oil. During May, the company produced 5,000 cases of lubricant. Each case contains 12 quarts of synthetic oil. To achieve this level of production, Slick purchased and used 16,500 gallons of direct materials at a cost of $20,472. It also incurred average direct labor costs of $14 per hour for the 3,884 hours worked in May by its production personnel. Manufacturing overhead for the month totaled $9,346, of which $2,200 was considered fixed. Slick's standard cost information for each case of synthetic motor oil is as follows.

Direct materials standard price $ 1.30 per gallon
Standard quantity allowed per case 3.25 gallons
Direct labor standard rate $ 16 per hour
Standard hours allowed per case 0.75 direct labor hours
Fixed overhead budgeted $ 2,600 per month
Normal level of production 5,200 cases per month
Variable overhead application rate $ 1.50 per case
Fixed overhead application rate ($2,600 ÷ 5,200 cases) 0.50 per case
Total overhead application rate $ 2.00 per case

Required:

a. Compute the materials price and quantity variances.

b. Compute the labor rate and efficiency variances.

c. Compute the manufacturing overhead spending and volume variances.

d. Prepare the journal entries to:

1. Charge materials (at standard) to Work in Process.

2. Charge direct labor (at standard) to Work in Process.

3. Charge manufacturing overhead (at standard) to Work in Process.

4. Transfer the cost of the 5,000 cases of synthetic motor oil produced in May to Finished Goods.

5. Close any over- or underapplied overhead to cost of goods sold.
Answer #1

ANSWER

(a):

Standard quantity of material for actual production = 5000*3.25 = 16,250 gallon

Actual quantity of material = 16,500 gallon

Standard price of material = $1.30

Actual price of material = $20,472 / 16,500 = $$1.2407

Material price variance = (SP - AP) * AQ = ($1.30 - $1.2407) * 16500 = $978 F

Material quantity variance = (SQ - AQ) * SR = (16,250 - 16,500) * $1.30 = $325 U

(b):

Standard hours of direct labor = 5,000 * 0.75 = 3750 hours

Standard rate of direct labor = $16 per hour

Actual hours of direct labor = 3,884

Actual rate of direct labor = $14 per hour

Direct labor rate variance = (SR - AR) * AH = ($16 - $14) * 3,884= $7,768 F

Direct labor efficiency variance = (SH - AH) * SR = (3750 - 3,884) * $16 = $2,144 U

(c):

Standard manufacturing overhead for actual production = (5000 * $1.50) + $2,600 = $10,100

Actual manufacturing overhead = $9,346

Manufacturing overhead spending variance = Budgeted manufacturing overhead - Actual manufacturing overhead

= $10,100 - $9,346= $754 F

Manufacturing overhead volume variance = Fixed overhead applied - Budgeted fixed overhead

= 5000 * $0.5 - $2,600 = $100 U

Solution (d)

Journal Entries - Slick Corporation
Event Particulars Debit Credit
1 Work In Process Dr (5000*3.25 * $1.30) $21,125.00
Direct material quantity variance Dr $325.00
        To Raw material inventory (16500*$1.30) $21,450.00
(To charge raw material on production)
2 Work In Process Dr (5000*0.75 * $16) $60,000.00
Direct labor efficiency variance Dr $2,144.00
        To Factory Wages (3,884*$16) $62,144.00
(To charge labor on production)
3 Work In Process Dr (5000*$2) $10,000.00
            To Manufactuirng overhead $10,000.00
(To charge MOH to production)
4 Finished goods inventory Dr $91,125.00
          To Work In Process $91,125.00
(To record transfer of completed units to finished goods)
5 Manufacturing overhead Dr $654.00
          To Cost of goods sold $654.00
(To close overapplied overhead to COGS)

===========================================

DEAR STUDENT,

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive ratings

*****************THANK YOU****************